# LOTUS: Enabling Semantic Queries with LLMs Over Tables of Unstructured and Structured Data

### v1

@@ Acknowledgements.
- mos ([n. d.]) [n. d.]. Mosaic AI Vector Search. https://docs.databricks.com
@@
- vec ([n. d.]) [n. d.]. Vector Database for Vector Search. https://www.pinecone.io/
- noa (2023a) 2023a. Large Language Models for sentiment analysis with Amazon Redshift ML (Preview) | AWS Big Data Blog. https://aws.amazon.com/blogs/big-data/large-language-models-for-sentiment-analysis-with-amazon-redshift-ml-preview/ Section: Amazon Redshift.
- noa (2023b) 2023b. Vespa - the big data serving engine. https://vespa.ai/
@@
- Andoni and Indyk (2008) Alexandr Andoni and Piotr Indyk. 2008. Near-optimal hashing algorithms for approximate nearest neighbor in high dimensions. Commun. ACM 51, 1 (Jan. 2008), 117–122. https://doi.org/10.1145/1327452.1327494
- Andoni et al. (2015) Alexandr Andoni, Piotr Indyk, Thijs Laarhoven, Ilya Razenshteyn, and Ludwig Schmidt. 2015. Practical and optimal LSH for angular distance. In Proceedings of the 28th International Conference on Neural Information Processing Systems - Volume 1 (NIPS’15). MIT Press, Cambridge, MA, USA, 1225–1233.
- Andoni and Razenshteyn (2015) Alexandr Andoni and Ilya Razenshteyn. 2015. Optimal Data-Dependent Hashing for Approximate Near Neighbors. In Proceedings of the forty-seventh annual ACM symposium on Theory of Computing (STOC ’15). Association for Computing Machinery, New York, NY, USA, 793–801. https://doi.org/10.1145/2746539.2746553
@@
- Baranchuk et al. (2018) Dmitry Baranchuk, Artem Babenko, and Yury Malkov. 2018. Revisiting the Inverted Indices for Billion-Scale Approximate Nearest Neighbors. https://doi.org/10.48550/arXiv.1802.02422 arXiv:1802.02422 [cs].
@@
- Ge et al. (2014) Tiezheng Ge, Kaiming He, Qifa Ke, and Jian Sun. 2014. Optimized Product Quantization. IEEE Transactions on Pattern Analysis and Machine Intelligence 36, 4 (April 2014), 744–755. https://doi.org/10.1109/TPAMI.2013.240 Conference Name: IEEE Transactions on Pattern Analysis and Machine Intelligence.
- Gionis et al. (1999) Aristides Gionis, Piotr Indyk, and Rajeev Motwani. 1999. Similarity Search in High Dimensions via Hashing. In Proceedings of the 25th International Conference on Very Large Data Bases (VLDB ’99). Morgan Kaufmann Publishers Inc., San Francisco, CA, USA, 518–529.
- Gong et al. (2020) Long Gong, Huayi Wang, Mitsunori Ogihara, and Jun Xu. 2020. iDEC: indexable distance estimating codes for approximate nearest neighbor search. Proceedings of the VLDB Endowment 13, 9 (May 2020), 1483–1497. https://doi.org/10.14778/3397230.3397243
@@
- Indyk and Motwani (1998) Piotr Indyk and Rajeev Motwani. 1998. Approximate nearest neighbors: towards removing the curse of dimensionality. In Proceedings of the thirtieth annual ACM symposium on Theory of computing (STOC ’98). Association for Computing Machinery, New York, NY, USA, 604–613. https://doi.org/10.1145/276698.276876
@@
- Jafari et al. (2020) Omid Jafari, Parth Nagarkar, and Jonathan Montaño. 2020. mmLSH: A Practical and Efficient Technique for Processing Approximate Nearest Neighbor Queries on Multimedia Data. In Similarity Search and Applications (Lecture Notes in Computer Science), Shin’ichi Satoh, Lucia Vadicamo, Arthur Zimek, Fabio Carrara, Ilaria Bartolini, Martin Aumüller, Björn Þór Jónsson, and Rasmus Pagh (Eds.). Springer International Publishing, Cham, 47–61. https://doi.org/10.1007/978-3-030-60936-8_4
@@
- Jégou et al. (2011) Herve Jégou, Matthijs Douze, and Cordelia Schmid. 2011. Product Quantization for Nearest Neighbor Search. IEEE Transactions on Pattern Analysis and Machine Intelligence 33, 1 (Jan. 2011), 117–128. https://doi.org/10.1109/TPAMI.2010.57 Conference Name: IEEE Transactions on Pattern Analysis and Machine Intelligence.
@@
- Li et al. (2020) Mingjie Li, Ying Zhang, Yifang Sun, Wei Wang, Ivor W. Tsang, and Xuemin Lin. 2020. I/O Efficient Approximate Nearest Neighbour Search based on Learned Functions. 2020 IEEE 36th International Conference on Data Engineering (ICDE) (April 2020), 289–300. https://doi.org/10.1109/ICDE48307.2020.00032 Conference Name: 2020 IEEE 36th International Conference on Data Engineering (ICDE) ISBN: 9781728129037 Place: Dallas, TX, USA Publisher: IEEE.
@@
- Liu et al. (2021) Wanqi Liu, Hanchen Wang, Ying Zhang, Wei Wang, Lu Qin, and Xuemin Lin. 2021. EI-LSH: An early-termination driven I/O efficient incremental c-approximate nearest neighbor search. The VLDB Journal 30, 2 (March 2021), 215–235. https://doi.org/10.1007/s00778-020-00635-4
- Lu and Kudo (2020) Kejing Lu and Mineichi Kudo. 2020. R2LSH: A Nearest Neighbor Search Scheme Based on Two-dimensional Projected Spaces. In 2020 IEEE 36th International Conference on Data Engineering (ICDE). 1045–1056. https://doi.org/10.1109/ICDE48307.2020.00095 ISSN: 2375-026X.
@@
- Malkov and Yashunin (2018) Yu A. Malkov and D. A. Yashunin. 2018. Efficient and robust approximate nearest neighbor search using Hierarchical Navigable Small World graphs. http://arxiv.org/abs/1603.09320 arXiv:1603.09320 [cs].
- Park et al. (2015) Yongjoo Park, Michael Cafarella, and Barzan Mozafari. 2015. Neighbor-sensitive hashing. Proceedings of the VLDB Endowment 9, 3 (Nov. 2015), 144–155. https://doi.org/10.14778/2850583.2850589
@@
- Sundaram et al. (2013) Narayanan Sundaram, Aizana Turmukhametova, Nadathur Satish, Todd Mostak, Piotr Indyk, Samuel Madden, and Pradeep Dubey. 2013. Streaming similarity search over one billion tweets using parallel locality-sensitive hashing. Proceedings of the VLDB Endowment 6, 14 (Sept. 2013), 1930–1941. https://doi.org/10.14778/2556549.2556574
@@
- Wang et al. (2021) Jianguo Wang, Xiaomeng Yi, Rentong Guo, Hai Jin, Peng Xu, Shengjun Li, Xiangyu Wang, Xiangzhou Guo, Chengming Li, Xiaohai Xu, Kun Yu, Yuxing Yuan, Yinghao Zou, Jiquan Long, Yudong Cai, Zhenxiang Li, Zhifeng Zhang, Yihua Mo, Jun Gu, Ruiyi Jiang, Yi Wei, and Charles Xie. 2021. Milvus: A Purpose-Built Vector Data Management System. In Proceedings of the 2021 International Conference on Management of Data (SIGMOD ’21). Association for Computing Machinery, New York, NY, USA, 2614–2627. https://doi.org/10.1145/3448016.3457550
@@
- Zhang et al. (2024) Peiyuan Zhang, Guangtao Zeng, Tianduo Wang, and Wei Lu. 2024. TinyLlama: An Open-Source Small Language Model. arXiv:2401.02385 [cs.CL] https://arxiv.org/abs/2401.02385
@@
- Zheng et al. (2020) Bolong Zheng, Xi Zhao, Lianggui Weng, Nguyen Quoc Viet Hung, Hang Liu, and Christian S. Jensen. 2020. PM-LSH: A fast and accurate LSH framework for high-dimensional approximate NN search. Proceedings of the VLDB Endowment 13, 5 (Jan. 2020), 643–655. https://doi.org/10.14778/3377369.3377374

### v2

@@ Acknowledgements.
- noa (2023) 2023. Large Language Models for sentiment analysis with Amazon Redshift ML (Preview) | AWS Big Data Blog. https://aws.amazon.com/blogs/big-data/large-language-models-for-sentiment-analysis-with-amazon-redshift-ml-preview/ Section: Amazon Redshift.
@@
- Dasgupta ([n. d.]) Sanjoy Dasgupta. [n. d.]. The hardness of k-means clustering. ([n. d.]).
@@
- Kang et al. (2021) Daniel Kang, John Guibas, Peter Bailis, Tatsunori Hashimoto, Yi Sun, and Matei Zaharia. 2021. Accelerating approximate aggregation queries with expensive predicates. Proceedings of the VLDB Endowment 14, 11 (July 2021), 2341–2354. https://doi.org/10.14778/3476249.3476285
@@
- Yaghmazadeh et al. (2017) Navid Yaghmazadeh, View Profile, Yuepeng Wang, View Profile, Isil Dillig, View Profile, Thomas Dillig, and View Profile. 2017. SQLizer: query synthesis from natural language. Proceedings of the ACM on Programming Languages 1, OOPSLA (Oct. 2017), 1–26. https://doi.org/10.1145/3133887 Publisher: Association for Computing Machinery.
- Yu et al. (2018) Tao Yu, Zifan Li, Zilin Zhang, Rui Zhang, and Dragomir Radev. 2018. TypeSQL: Knowledge-based Type-Aware Neural Text-to-SQL Generation. https://doi.org/10.48550/arXiv.1804.09769 arXiv:1804.09769 [cs].
@@
- Zelle and Mooney (1996) John M Zelle and Raymond J Mooney. 1996. 1996-Learning to Parse Database Queries Using Inductive Logic Programming. (1996).
- Zhang et al. (2024) Bin Zhang, Yuxiao Ye, Guoqing Du, Xiaoru Hu, Zhishuai Li, Sun Yang, Chi Harold Liu, Rui Zhao, Ziyue Li, and Hangyu Mao. 2024. Benchmarking the Text-to-SQL Capability of Large Language Models: A Comprehensive Evaluation. https://doi.org/10.48550/arXiv.2403.02951 arXiv:2403.02951 [cs].